Aug 12, 2010 Readout

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET and EVO session
  • Phone in CC L207 is 757-269-7084
  • Phone:
    • +1-800-377-8846 : US
    • +1-888-276-7715 : Canada
    • +1-302-709-8424 : International
    • then enter participant code: 77438230# (remember the "#"). Phone connection only upon request.

Items for followup from previous meeting

  1. Testing goals and setup (Serguei)
  2. Will to find out if there is any feedback from USM on terms and conditions from our procurement package.
  3. Will to update statement of work for USM contract and develop detailed schedule.
  4. Collect this information of testing on the wiki for ease of presenting information. (USM)

Documents to Review

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. Bcal thresholds and resolution at low energies (PDF Slides) (Matt)
  4. SiPM Radiation Hardness Test (Update from Yi)
  5. Geometry
  6. Update - USM
    • Status of statement of work (SOW)
  7. Simulation runs
  8. Cosmic-ray setup at UofR
  9. Light guides


Attendees: Eugene, Tim, Fernando, Beni, Yi, Chuck, John (JLab), Andrei, Zisis (Regina)

  1. Announcements: None
  2. Review of Action Items: USM did not connect, so this discussion was deferred to a future meeting.
  3. Bcal thresholds and resolution at low energies: Eugene summarized Matt's work in that his numbers and Elton's are consistent and the lesser radiation load for the upstream SiPMs does not confer a significant improvement (only 10-15%). Regina will also have a close look at the numbers.
  4. SiPM Radiation Hardness Test Yi provided an update of the planned tests at JLab, shown on his Wiki page.
    • The DAQ tests aim at studying the ADC spectrum in order to extract the Dark rate and gain information. The additional radiation of 30 rem corresponds roughly to 3 years of running in Hall D. The AmBe tests will be done on the SiPMs without bias on them.
  • Tagger 2x2mm^2 devices from Photonique will also be tested.
  • RadCon is busy with calibrations, so those tests may start in two weeks from now and will probably need a month to be completed, although preliminary results will be made known as they become available.
  • 12 new 1x1mm^2 SiPMs will be ordered from Hamamatsu for these tests. Fernando will provide the contact person at Hamamatsu, and said that these are off-the-shelf, catalog items so expedited shipment can be pursued.
  1. Geometry:
    • A brief continuation of the light guide (LG) geometry and mounting took place, along the lines of the previous meeting. The options include mounting the 40 LGs (as a single unit or glued to a Lucite plate) on each face at the "bench" and then inserting the modules into the magnet, or inserting the modules first and then mounting the LGs in situ. Horizontal gluing may require UV fast-curing epoxy. If we opt for FM devices with the long LGs, the straight section can be glued on the bench and the curved section in situ, after the modules have been mounted inside the solenoid.
    • KLOE mounted the light guides with the modules oriented vertically but they also implemented 160 um Aluminum tape, at several spots along their module's length. Use of such tape has implications for our O-rings.
  2. Update - USM: no report.
    • Status of statement of work (SOW)
  3. Simulation runs: Irina's simulations of the full cosmics setup and cosmic ray distributions were completed, and are being used by Andrei in his cosmics analysis.
  4. Cosmic-ray setup at UofR: a five-point length scan of the module was completed, and now the trigger counters are in the middle position, for cross-check measurements. A report will be presented at the September Collaboration Meeting.
  5. Light guides: no report.